Using synthetic subtraction,
Write the polynomial equation of degree 4 with leading
coefficient 1
that has roots at ­2, ­1, 3, and 4.

Answer by Edwin McCravy(20086)

Using synthetic subtraction,
Write the polynomial equation of degree 4 with leading
coefficient 1
that has roots at ­2, ­1, 3, and 4.
I don't know what "synthetic subtraction" means.  But here's
the normal way people do your problem.

  x=2;   x=1;   x=3;   x=4;
x-2=0; x-1=0  x-3=0  x-4=0

(x-2)(x-1)(x-3)(x-4) = 0

Multiply that all the way out and get

x4-10x3+35x2-50x+24 = 0

Edwin